Suppose the consumption function of a society is:C=600+0.85Y where C=Total Consumption;Y=Personal disposable income.Identify the following:1)Autonomous consumption.2)Marginal propensity to consume;3)When Y=1000,what will be the induced consumption?4)If income is increased to 3000,what will happen to the total consumption?5)What is the autonomous consumption when income is 1000?
1
Expert's answer
2015-03-10T10:37:09-0400
C=600+0.85Y where C=Total Consumption;Y=Personal disposable income.
As C = Ca + c*Y:
1) Autonomous consumption is Ca = 600.
2) Marginal propensity to consume is c = 0.85;
3) When Y=1000, the induced consumption is 0.85*1000 = 850
4) If income is increased to 3000, the total consumption is C = 600 + 0.85*3000 = 3150
5) When income is 1000, the autonomous consumption is still 600.